# Countering adversarial evasion in regression analysis

### v1

@@ 2 Training resilient prediction functions
Now, we can define the block functions, $\mathcal{G}$ and $\mathcal{H}$ as below,
$\mathcal{G}(z):=\begin{bmatrix}g(X)\\ g(X)\\ 0\end{bmatrix},\qquad\mathcal{H}(z,\xi;B_{i}):=\begin{bmatrix}\nabla_{w}L^{\textup{p}}_{w}(z,\xi)\\ \nabla_{X}L^{\textup{p}}_{X}(z,\xi)\\ \nabla_{X}\ell^{\textup{p}}(z,\xi)\end{bmatrix}$.
Then, as outlined in the solution method in [37, Section 3], we can find stationary points of the bilevel program by solving the following system,
$\zeta\geq 0,\,\mathcal{G}(z,\zeta)\leq 0,\,\zeta^{T}\mathcal{G}(z,\zeta)=0,\,\mathcal{H}(z,\zeta;B_{i})=0.$
This system can be converted into a system of equalities via the Fischer-Burmeister function,
$\Phi(z,\zeta;B_{i})=\begin{pmatrix}\mathcal{H}(z,\zeta;B_{i})\\ \varphi_{\textup{FB}}\left(\zeta,-\mathcal{G}(z,\zeta)\right)\end{pmatrix}=0,$ (9)
where $\xi:=(z,\zeta)\in\mathbb{R}^{p}\times\mathbb{R}^{m\times p}\times\mathbb{R}^{2b+1}$ is a block variable and $\varphi_{\textup{FB}}$ is the Fischer-Burmeister, defined as
$\varphi_{\textup{FB}}(a,b)=\sqrt{a^{2}+b^{2}}-a-b.$
This system can then be solved using the Levenberg-Marquardt algorithm given in [37, Algorithm 1], where $\Psi:(\mathbb{R}^{p}\times\mathbb{R}^{mp})\times\mathbb{R}^{2b+1}$ is defined as,
[…]
The resulting weights of this solution can be used to construct a prediction function that has accounted for adversarial influence during the training process. With the inclusion of the constraints in (6), we can be confident that the adversary’s data retains the intended nature and realism after the transformation.
@@ 3 Numerical experiments
We assess the resilience of the bilevel model to strategic adversarial modifications. To do so, we design experiments on two datasets, the first of which, named Wine Quality records 11 features of 4898 bottles of red wine such as acidity and alcohol content. A corresponding quality score between 0100-10 is measured and recorded for each bottle. For this dataset, we consider scenarios where wine producers might lie when reporting the features of the wine or perhaps bribe quality assessors to alter the results in order to appear to be producing wine of a higher quality than they actually are. The second dataset, named Real Estate, contains the values of 6 features of 414 houses such as their age and distance to the nearest train station along with the corresponding sale price. We consider scenarios where sellers might lie about features of the house in order to be able to charge a higher price than deserved.
Further to this, we use the numerical experiments in this section as an opportunity to explore a particularly useful aspect of the constrained model which was not exploitable in [37]. Specifically, we demonstrate how, unlike existing pessimistic bilevel approaches to adversarial training, our models allow us to investigate which features of the adversary’s data were changed and hence give us insight into the most vulnerable aspects of the prediction models. In the current proposal for a pessimistic bilevel approach to evasion attacks, [5], pass the adversary’s data through a feature map. This plays a vital role in the proposed solution process, as the resulting feature space can be assumed to be unrestricted, which is convex. This, coupled with a strictly convex objective function, grantees a unique solution to the adversary’s problem in the lower-level, allowing the model to be reorganised as a pessimistic bilevel optimisation problem, which is easier to solve. However, since our model makes no assumptions on the convexity of the adversary’s problem or the uniqueness of its solution, we can leave the adversary’s data in its original feature space. Therefore, we can observe the adversary’s solution and infer the extent to which features must be changed in order to trick the learner. This analysis could not be performed in the experiments in [37], as the text-based data needed to be embedded before analysis. This embedding is not bijective and, consequently, the value of the adversary’s data at the solution provided no useful insight into the adversary’s strategy. However, in this chapter, we take the opportunity to test the bilevel model on some datasets that require no feature manipulation and hence allow us to investigate the adversary’s movements.
@@
$X^{\text{test}}_{i}:=\operatorname*{argmin}_{x\in\mathbb{R}^{q}}\left\{\ell(\sigma(w,x),Z^{\text{test}}_{i})\,\mid\,g(x,x^{0};\delta^{\text{test}}_{i})\leq 0\right\},\quad i\in I,$
where $x^{0}\in\mathbb{R}^{q}$ is the initial values of $x$. We set $T$ to be $10\%$ of the size of the test set. All features and labels are normalised to the range $(0,1)$ and we set the adversarial perturbation to be $\Delta=2\text{std}(Y^{\text{test}})$ where $\text{std}(Y^{\text{test}})$ is the standard deviation of the test labels. We compare our model to that of a typically trained linear regression predictor, named LinReg, for various values of $m$ and plot in Figure 1. We can see for both datasets a similar trend of an initially high mean square error (MSE) for $m=1$, before a drastic decrease to an optimal value of $m=2$. Following this, we see a gradual increase in MSE as $m$ increases, before plateauing around an MSE roughly equal to that of the traditional predictor.
@@
Under this measure, a larger distance implies that the adversary needed to transform the data by a larger extent in order to sufficiently fool the leaner’s prediciton model. We plot the average movements for each feature and for each dataset in Figure 2. For the Wine quality, we see the most movement in alcohol followed closely by volatile acidity and residual sugar. These results give the insight into how much an adversary must modify each feature in order to fool the learner. It is clear, for example, that the alcohol content must be modified considerably more than the citric acid to generate the most effective wine. On the other hand, the citric acid content required will little change to achieve a value sufficient to fool the learner’s predictor. The citric acid, therefore, might be considered a particularly vulnerable feature that could easily be exploited and perhaps should be treated with more scrutiny. In the real estate data, we see the distance to the neaest MRT station as the most resilient feature, requiring relatively substantial change to effect the learner’s predictor. Meanwhile, the transaction data and longitude appear to be the easiest to exploit.
@@
It is clear from the experimental results that the constrained pessimistic bilevel model provides an effective method of accounting for adversarial manipulation during the training process of regression scenarios. The lower-level successfully anticipates the movements of an adversary, which leads to improved performance by the prediction model trained in the upper-level. We see a significant change in performance as we increase the number of instances available to be manipulated by the adversary. In particular, we note that a balance needs to be struck between allowing the adversary enough instances that they have sufficient influence over then training process, while not too many that the model overestimates the influence of the adversary. Finally, we demonstrated how, unlike the work in [37], we could investigate the values of the adversary’s data at the solution. This allowed us to identify which features were most vulnerable to exploitation.
@@ 4 Conclusion
Throughout this paper, we demonstrated how the pessimistic bilevel model proposed by [37] can be adapted to adversarial regression scenarios. In particular, by anticipating adversarial movement in the lower-level, we can train prediction functions that are resilient to adversaries who attempt to influence particular outcomes from a prediction model by strategically transforming their data. For example, in the context of property valuation, an adversary might purposefully provide falsified information to be achieve higher valuations and hence charge a higher selling price. While there exist pessimistic bilevel approaches to adversarial evasion scenarios, these models exploit strong assumptions about the convexity of the adversary’s problem and the uniqueness of their solution to reformulate the program into its optimistic variant, which, while easier to solve, oversimplifies the capabilities of the adversary. Since our model makes no such assumptions, we retain the pessimistic aspect of the bilevel model, allowing us to accurately capture the antagonistic nature of these scenarios.
To assess the again simulated adversarial influence on a number of applications, including insurance appraisal, wine quality appraisal and real estate pricing. Throughout numerical experiments, we demonstrated the ability of this model to train resilient predictors that provided improved performance over traditionally-trained predictors. Further to this, we again investigated varying the number of samples available to the adversary and the similarity threshold and observed a clear pattern in the form of a trade-off between allowing the adversary enough freedom and influence to generate sufficient movement that impacts the training process, while not too much such that the adversary provides too much influence over the training process, leading to overly pessimistic predictors.
@@ Appendix A Derivatives for the leader and follower
Let $x\in\mathbb{R}^{q}$ be a sample of data with corresponding label $y\in\mathbb{R}$ abd let $w\in\mathbb{R}^{q}$ be the weights of the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\sigma(w,x)=w^{T}x.$
We define the upper-level (learner’s) loss function $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\mathcal{L}(\sigma(w,x),y):=\|\sigma(w,x)-y\|_{2}^{2}.$
Let $z\in\mathbb{R}$ be the adversary’s target label. We define the lower-level (adversary’s) loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as,
$\ell(\sigma(w,x),y):=\|\sigma(w^{T}x)-z\|_{2}^{2}.$
Let $D\in\mathbb{R}^{mq}$ be the static data with corresponding labels $\gamma\in\mathbb{R}^{n}$ and let $X\in\mathbb{R}^{mq}$ be the adversary’s data with corresponding labels $Y\in\mathbb{R}^{m}$. The upper-level objective $F:\mathbb{R}^{q}\times\mathbb{R}^{mq}\rightarrow\mathbb{R}$ is defined as,
$F(w,X):=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}(\sigma(w,D_{i}),\gamma_{i})+\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}(\sigma(w,X_{i}),Y_{i})+\frac{1}{\rho}\|w\|_{2}^{2}.$
The derivative of the upper-level (leader’s) loss with respect to the learner’s weights is given by,
$\frac{\partial F}{\partial w}(w,X)=\frac{2}{n}D^{T}(w^{T}D-\gamma)+\frac{2}{m}X^{T}(w^{T}X-Y)+\frac{2}{\rho}w.$
@@
$\frac{\partial^{2}F}{\partial ww}(w,X)=\frac{2}{n}D^{T}D+\frac{2}{m}X^{T}X+\frac{2}{\rho}$
The derivative of the upper-level loss function with respect to the adversary’s data is given by
@@
$\frac{\partial^{2}\mathcal{L}}{\partial X_{ij}X_{kl}}=\begin{cases}\frac{2}{m}w_{j}w_{l}&i=k\\ 0&\text{otherwise}\end{cases}.$
Let $X^{0}\in\mathbb{R}^{mq}$ be the start point of the adversary’s data. We define the constraints $g:\mathbb{R}^{mq}\rightarrow(-1,1)$ of the lower-level component-wise as
$g_{i}(X)=\delta-d(X_{i},X_{i}^{0}),\;i\in\{1,\dots,m\},$
where
$d(X_{i},X_{i}^{0})=\frac{X_{i}\cdot X_{i}^{0}}{\|X_{i}\|\|X_{i}^{0}\|},\;i\in\{1,\dots,m\},$
and $\delta\in\mathbb{R}$ is the similarity threshold. The derivative of the constraints with respect to the classifier weights is 0. The derivative with respect to the adversary’s data is given by the following cases,
$\frac{\partial g_{i}(X)}{\partial X_{jk}}=\begin{cases}\frac{X_{ik}^{0}}{\|X_{i}\|\cdot\|X_{i}^{0}\|}-d(X_{i},X_{i}^{0})\frac{X_{ik}}{\|X_{i}\|^{2}}&i=j\\ 0&i\neq j\end{cases}.$
The second derivative with respect to the adversary’s data is given by the cases,
$\frac{\partial g_{i}(X)}{X_{jk}X_{lc}}=\begin{cases}\frac{X_{ic}X_{ik}^{0}+X_{ik}X_{ic}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}X_{ic}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}&i=j=l,k\neq c\\ \frac{2X_{ik}X_{ik}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}^{2}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}+\frac{d(X_{i},X_{i}^{0})}{\|X_{i}\|^{2}}&i=j=l,k=c\\ 0&\text{otherwise}.\end{cases}$

### v2

@@ 2 Training resilient prediction functions
Based on this notation, the system described by (12a)–(12e) can then be restated as
$\zeta\geq 0,\;\;\mathcal{G}(z)\leq 0,\;\;\zeta^{T}\mathcal{G}(z)=0,\;\;\mathcal{H}(z,\xi)=0,$ (13)
which can equivalently be written as the following system of equations
$\left\{\begin{array}[]{l}\mathcal{H}(z,\zeta)=0,\\ \vartheta_{\textup{FB}}\left(\zeta_{i},-\mathcal{G}_{i}(z,\zeta)\right)=0,\;i=1,\ldots,2m.\end{array}\right$. (14)
Here, $\vartheta_{\textup{FB}}$ corresponds to the so-called Fischer-Burmeister function [13], which is defined by
$\vartheta_{\text{FB}}(a,b):=\sqrt{a^{2}+b^{2}}-(a+b)\;\mbox{ for }\;(a,b)\in\mathbb{R}^{2}.$
Thanks to this function, the second equation in (13) is equivalent to the complementarity conditions in (12d)–(12e), written in compact form in (13), while considering them in pairs.
Clearly, the stationary conditions (12a)–(12e) of problem (9) - (10) have been written as a system of equations in (14). This system is overdetermined, with $m$ more equations than variables, making the Levenberg-Marquardt method, specifically, the global nonsmooth Levenberg–Marquardt method for mixed nonlinear complementarity systems developed in [17], a suitable choice to solve it. A version of this algorithm which has been appropriately modified for the pessimistic adversarial bilevel program by introducing additional stopping criteria, is given in [27, Algorithm 2]. We apply this same algorithm here to solve system (14), using the derivative formula provided in Appendix A, and extract the solution $z^{*}=(w^{*},X^{*})$. This solution comprises the weights, $w^{*}$, of a prediction function that has accounted for adversarial influence during the training process as well as the corresponding transformed adversary’s data, $X^{*}$. Unlike the framework proposed in the previous work in [27], it is not required that the adversary’s data is embedded ahead of analysis. Consequently, we can use $X^{*}$ to observe the value of the adversary’s solution in its original feature space and investigate the nature of their transformation.
@@ 3 Numerical experiments
We assess the resilience of the bilevel model to strategic adversarial modifications. To do so, we design experiments on two datasets, the first of which, named Wine Quality records 11 features of 4898 bottles of red wine such as acidity and alcohol content. A corresponding quality score between 1101-10 is measured and recorded for each bottle. For this dataset, we consider scenarios where wine producers might lie when reporting the features of the wine or perhaps bribe quality assessors to alter the results in order to appear to be producing wine of a higher quality than they actually are. The second dataset, named Real Estate, contains the values of 6 features of 414 houses such as their age and distance to the nearest train station along with the corresponding sale price. We consider scenarios where sellers might lie about features of the house in order to be able to charge a higher price than deserved. Further to this, we use the numerical experiments in this section to explore a particularly useful aspect of the constrained model which, due to the requirement of feature transformations, was not exploitable in previous works. Specifically, we demonstrate how, unlike existing pessimistic bilevel approaches to adversarial training, our models allow us to investigate which features of the adversary’s data were changed and hence give us insight into the most vulnerable aspects of the prediction models.
@@
$X^{\text{test}}_{i}:=\operatorname*{argmin}_{x\in\mathbb{R}^{q}}\left\{\ell\left(\sigma(w^{\text{init}},x\right),Z^{\text{test}}_{i})\,\left|\,\delta^{\text{test}}_{i}-\frac{x\cdot Z^{\text{test}}_{i}}{\|x\|\|Z^{\text{test}}_{i}\|}\leq 0\right.\right\},\;\forall i\in I.$
where $x^{0}\in\mathbb{R}^{q}$ is the initial values of $x$. We set $T$ to be $10\%$ of the size of the test set. All features and labels are normalised to the range $(0,1)$ and we set the adversarial perturbation to be $\Delta=2\text{std}(Y^{\text{test}})$ where $\text{std}(Y^{\text{test}})$ is the standard deviation of the test labels.
We compare our model to that of a typically trained linear regression predictor, named LinReg. While the pessimistic model in [4] was intended for use on classification tasks, its prediction function and loss functions can be easily substituted for linear regression variants which still satisfy their assumptions about strict convexity. We name this B&S and use it as a comparison to the existing pessimistic approach to adversarial regression analysis. The mean square error for these models as well as our model for various values of the adversary’s sample size, $m$, and the similarity threshold, $\delta$, are plotted in Figure 1. We can see for both datasets a similar trend of an initially high mean square error (MSE) for low values of $m$, before a decrease to an optimal value. Following this, we observe an increase in MSE, before plateauing around an MSE roughly equal to that of the traditional predictor. Both experiments see an optimal MSE when $\delta=0.95$ and when $m$ is fairly low. For the wine dataset, the optimal value falls at $m=1$, while on the real estate dataset, we see the optimal value fall at $m=2$. Although, we note that the real estate dataset also sees similarly good performance for $m$ in the range of 132313-23. In general, we observe a general pattern in that the adversary should be granted enough freedom and influence over the training process to capture the nature of adversarial attacks, while not too much influence that the model becomes over-estimates the power of the adversary and hence suffers in performance.
@@
Under this measure, a larger distance implies that the adversary needed to transform the data by a larger extent in order to sufficiently fool the leaner’s prediction model. We plot the average movements for each feature and for each dataset in Figure 2. For the Wine quality, we see the most movement in residual sugar followed by PH and sulfur dioxide. These results give the insight into how much an adversary must modify each feature in order to fool the learner. It is clear, for example, that the alcohol content must be modified considerably to generate the most effective wine. While chlorides, on the other hand, required little change to achieve a value sufficient to fool the learner’s predictor. Chlorides, therefore, might be considered a particularly vulnerable feature that could easily be exploited and perhaps should be treated with more scrutiny. In the real estate data, we see the distance to the nearest MRT station as the most resilient feature, requiring relatively substantial change to affect the learner’s predictor. Meanwhile, the transaction data and longitude appear to be the easiest to exploit.
@@
It is clear from the experimental results that the pessimistic bilevel model with constrained adversarial movement provides an effective method of accounting for adversarial manipulation during the training process of regression scenarios. The lower-level successfully anticipates the movements of an adversary, which leads to improved performance by the prediction model trained in the upper-level. We see a significant change in performance as we increase the number of instances available to be manipulated by the adversary. In particular, we note that a balance needs to be struck between allowing the adversary enough instances that they have sufficient influence over then training process, while not too many that the model overestimates the influence of the adversary. Finally, we demonstrated how, unlike existing works, we could investigate the values of the adversary’s data at the solution. This allowed us to identify which features were most vulnerable to exploitation.
@@ 4 Conclusion
In this article, we proposed a pessimistic bilevel program to model adversarial evasion in regression scenarios. In particular, by anticipating adversarial movement in the lower-level, we can train prediction functions that are resilient to adversaries who attempt to influence particular outcomes from a prediction model by strategically transforming their data. For example, in the context of property valuation, an adversary might purposefully provide falsified information to achieve higher valuations and hence charge a higher selling price. While there exist pessimistic bilevel approaches to adversarial evasion scenarios, these models exploit strong assumptions about the convexity of the adversary’s problem and the uniqueness of their solution to reformulate the program into its optimistic variant, which, while easier to solve, oversimplifies the capabilities of the adversary. Since our model makes no such assumptions, we retain the pessimistic aspect of the bilevel model, allowing us to accurately capture the antagonistic nature of these scenarios. Furthermore, with the introduction of lower-level constraints, which restrict the adversary’s movement, preventing nonsensical transformations and ensuring the adversary’s data remains plausibly realistic, leading to more accurate prediction functions.
We simulated adversarial influence to create test sets for wine quality appraisal and real estate pricing. Numerical experiments demonstrated the ability of the bilevel model to train resilient predictors that provided improved performance over existing methods. Further to this, we investigated varying the number of samples available to the adversary and observed a clear pattern in the form of a trade-off between allowing the adversary enough freedom and influence to generate sufficient movement that impacts the training process, while not too much such that the adversary provides too much influence over the training process, leading to overly pessimistic predictors. Additionally, we investigated a feature of the pessimistic bilevel model which previous works were not able to exploit. Specifically, since we do not transform the adversary’s data before analysis, we were able to observe the adversary’s solution in its original feature space and measure the extent to which the adversary modified each feature. From this we could identify which features were most vulnerable to attacks.
@@ Appendix A Derivatives for the leader and follower
Let $x\in\mathbb{R}^{q}$ be a sample of data with corresponding label $y\in\mathbb{R}$ and let $w\in\mathbb{R}^{q}$ be the weights of the learner’s prediction function, $\sigma:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ defined as in (1). We define the upper-level (learner’s) loss function $\mathcal{L}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ as it is in (2). Let $z\in\mathbb{R}$ be the adversary’s target label.
Let $D\in\mathbb{R}^{mq}$ be the static data with corresponding labels $\gamma\in\mathbb{R}^{n}$ and let $X\in\mathbb{R}^{mq}$ be the adversary’s data with corresponding labels $Y\in\mathbb{R}^{m}$. The upper-level objective $F:\mathbb{R}^{q}\times\mathbb{R}^{mq}\rightarrow\mathbb{R}$ is defined as it is in (4). The derivative of the upper-level (leader’s) loss with respect to the learner’s weights is given by,
$\nabla_{w}F(w,X)=\frac{2}{n}D^{T}(w^{T}D-\gamma)+\frac{2}{m}X^{T}(w^{T}X-Y)+\frac{2}{\rho}w.$
@@
$\nabla^{2}_{ww}F(w,X)=\frac{2}{n}D^{T}D+\frac{2}{m}X^{T}X+\frac{2}{\rho}$
The derivative of the upper-level objective function with respect to the adversary’s data is given by
$\frac{\partial F}{\partial X_{ij}}(w,X):=\frac{1}{m}\sum_{k=1}^{m}\frac{\partial\mathcal{L}}{\partial X_{ij}}(\sigma(w,X_{k}),Y_{k})=\frac{1}{m}\frac{\partial\mathcal{L}}{\partial X_{ij}}(\sigma(w,X_{i}),Y_{i}),$
where
@@
$\frac{\partial^{2}\mathcal{L}}{\partial X_{ij}\partial X_{kl}}=\begin{cases}\frac{2}{m}w_{j}w_{l}&i=k\\ 0&\text{otherwise}.\end{cases}$
The derivative of the upper-level objective function with respect to the adversary’s data is given by
$\frac{\partial f}{\partial X_{ij}}(w,X)=\frac{1}{m}\sum_{i=k}^{m}\frac{\partial\ell}{\partial X_{ij}}(\sigma(w,X_{k}),Y_{k})=\frac{1}{m}\frac{\partial\ell}{\partial X_{ij}}(\sigma(w,X_{k}),Y_{k}).$
where the lower-level (adversary’s) loss function $\ell:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ is defined as it is in (3). Note that the adversary’s loss function can be expressed in terms of the leader’s loss function,
$\ell(\sigma(w,x),z)=\mathcal{L}(\sigma(w,x),z),$
for some weights $w\in\mathbb{R}^{q}$, data $x\in\mathbb{R}^{q}$ and target label $z\in\mathbb{R}$. Therefore, we can express the derivative of the lower-level objective function with respect to the adversary’s data as follows
$\frac{\partial f}{\partial X_{ij}}(w,X)=\frac{1}{m}\frac{\partial\mathcal{L}}{\partial X_{ij}}(\sigma(w,X_{i}),Z_{i})\quad i=1,\dots,m,\;j=1,\dots,q.$
The second derivative with respect to the adversary’s data is then given as
$\frac{\partial^{2}f}{\partial X_{ij}\partial X_{kc}}(w,X)=\frac{1}{m}\frac{\partial^{2}\mathcal{L}}{\partial X_{ij}\partial X_{jk}}(\sigma(w,X_{i}),Z_{i})\quad i,k=1,\dots,m,\;j,c=1,\dots,q.$
Finally, the derivative with respect to the learner’s weights and the adversary’s data is given by
$\frac{\partial^{2}f}{\partial w_{i}\partial X_{jk}}(w,X)=\frac{1}{m}\frac{\partial^{2}\mathcal{L}}{\partial w_{i}\partial X_{jk}}(\sigma(w,X_{i}),Z_{i})\quad i,j=1,\dots,m,\;k=1,\dots,q.$
Let $X^{0}\in\mathbb{R}^{mq}$ be the start point of the adversary’s data and let the lower-level constraints $g:\mathbb{R}^{mq}\rightarrow(-1,1)$ be defined as in (6), where each constraint function $g_{i}(X):\mathbb{R}^{q}\rightarrow\mathbb{R},\;i=1,\dots,m$ measures the cosine similarity between the adversary’s data and its original position, as given by (8), where $\delta\in\mathbb{R}$ is the similarity threshold. The derivative of the constraints with respect to the classifier weights is 0. The derivative with respect to the adversary’s data is obtained as
$\frac{\partial g_{i}(X)}{\partial X_{jk}}=\begin{cases}\frac{X_{ik}^{0}}{\|X_{i}\|\cdot\|X_{i}^{0}\|}-d(X_{i},X_{i}^{0})\frac{X_{ik}}{\|X_{i}\|^{2}}&i=j,\\ 0&i\neq j.\end{cases}$
The second derivative with respect to the adversary’s data is given by the cases
$\frac{\partial^{2}g_{i}(X)}{\partial X_{jk}\partial X_{lc}}=\begin{cases}\frac{X_{ic}X_{ik}^{0}+X_{ik}X_{ic}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}X_{ic}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}&i=j=l,\;\,k\neq c,\\ \\ \frac{2X_{ik}X_{ik}^{0}}{\|X_{i}\|^{3}\|X_{i}^{0}\|}-\frac{3X_{ik}^{2}d(X_{i},X_{i}^{0})}{\|X_{i}\|^{4}}+\frac{d(X_{i},X_{i}^{0})}{\|X_{i}\|^{2}}&i=j=l,\;\,k=c,\\ \\ 0&\text{otherwise}.\end{cases}$